OpenStudy (anonymous):

solve using substitution method 4m+n=18 m-7n=19

OpenStudy (anonymous):

28+7n=126 m-7n=19

OpenStudy (anonymous):

29M=145 M=5

OpenStudy (anonymous):

29m=145 m=145/29 m=5

OpenStudy (anonymous):

N=-2

OpenStudy (anonymous):

n=2

OpenStudy (anonymous):

n=-2

OpenStudy (anonymous):

those answers are wrong

OpenStudy (anonymous):

well just take the second equation and do this.... m-7(4m+n)=19

OpenStudy (anonymous):

m-28m+7n=19

OpenStudy (anonymous):

-27m+7n=19

OpenStudy (anonymous):

4m+n=18 m-7n=19 therefore m=7n +19 substitute into 4m+n=18 4(7n +19) = 18 28n -76 =18 subtract 76 from both sides 28n = -58 divide by 28 both sides n =29/14 Now feed this value of n into the other equation to find m

OpenStudy (anonymous):

Sorry thatshould read -29/14

OpenStudy (anonymous):

can you help me do that?

OpenStudy (anonymous):

Yes 4m+n=18 but n =-29/14 So 4m - (-29/14) = 18 4m +29/14 = 18 4m = 18 -29/14 4m = 223/14 divide by 4 m =223/56 Now check both values of m and n in the second equation.

OpenStudy (anonymous):

First take the first equation and solve for N. now take what you got for N and plug it into the second equation for the N and solve for M. then you should have the two answers.

OpenStudy (anonymous):

shouldn't it have been 4(7n+19)+n= 18?

OpenStudy (anonymous):

Chris, I understand it enough to know that is part of solving. I really do write down these steps and study them...If I continue getting the wrong answers I put the questions on here.

OpenStudy (anonymous):

Gianfranco which one gets listed first the m or n? I know in x and y x always goes first.

OpenStudy (anonymous):

M=5 n = -2 sorry rachel it is very early here!

OpenStudy (anonymous):

lol it is ok. so M goes first?

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

go figure

OpenStudy (anonymous):

Hav you graphed the lines to see the point of intersection as that will also give you a rtesult?

OpenStudy (anonymous):

Should it go (m,n)?

OpenStudy (anonymous):

I only have to solve for these specific problems but yes I did do some of those problems

OpenStudy (anonymous):

thanks for the medal shad1

OpenStudy (anonymous):

id give you one but i dont know how

OpenStudy (anonymous):

4m+n=18 m-7n=19 so m=19+7n so 4(19+7n)+n=18 29n=18-76=-58 so n=-2 so m=5